OpenStudy (anonymous):

I have six terms: 5,7,11,19,35,67, what is a1

OpenStudy (solomonzelman):

What do you think?

OpenStudy (solomonzelman):

\[a1~~~~~usually~~written~~as~~a_1~~~is~~the~~first~~term~~of~~the~~sequence.\]

OpenStudy (anonymous):

a(some number goes here) represents a numbered place in the series

OpenStudy (solomonzelman):

\[A_2=A_1+(1\times 2)\]\[A_3=A_2+(2\times 2)\]\[A_4=A_3+(3\times 2)\]\[A_5=A_4+(4 \times 2)\]Do you see the pattern?